a plumber has a 20 foot piece of pvc pipe. how many 7/5 foot pieces can be cut from the 20 foot piece?
Answer by addingup(3677)

(20/1)/(7/5) = 20/1 x 5/7 = (20 x 5)/(1 x 7) = 100/7 = 14 pieces and you have a small remnant.